Wednesday, 19 August

Plenary Lectures

9:30 – 10:30  Auditorium Maximum H 105

Peter W. Shor, AT&T Labs, Florham Park, USA Quantum Computing (I: p.19)

Chair: Volker Strassen, Universität Konstanz, Konstanz, Germany

11:00 – 12:00  Auditorium Maximum H 105

Ehud Hrushovski, Hebrew University of Jerusalem, Jerusalem, Israel Geometric Model Theory (I: p.9)

Chair: Ronald Jensen, Humboldt–Universität, Berlin, Germany

ICM’98 – Program  1